OpenStudy (princessaurora):

*Will fan and Medal* Of the 16 cards drawn from a standard deck of 52 cards, 5 of the cards are spades. Based on the cards that were drawn, the experimental probability of drawing a spade is BLANK , and the difference of the experimental and theoretical probabilities of drawing a spade from a standard deck of cards is BLANK. What are the blanks?

OpenStudy (pawanyadav):

Experimental probability=5/16 And theoretical probability= (favourable outcome)/total outcome =13/52=1/4

OpenStudy (princessaurora):

Thank you a million times!! :)

OpenStudy (pawanyadav):

You are welcome